English

A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of 5 successes.

Advertisements
Advertisements

Question

A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of 5 successes. 

Sum
Advertisements

Solution

Let X = number of successes, i.e., number of odd numbers.

p = probability of getting an odd number in a single throw of a die

∴ p = `3/6 = 1/2` and

q = `1 - "p" = 1 - 1/2 = 1/2`

Given: n = 6

∴ X ∼ B`(6, 1/2)`

The p.m.f. of X is given by

`p("X = x") = "^nC_x  p^x  q^(n - x)`

i.e. p(x) = `"^6C_x (1/2)^x (1/2)^(6 - x)`

` = "^6C_x (1/2)^6,` x = 0, 1, 2, ...,6

P(5 successes) = P[X = 5]

`= p(5) = "^6C_5 (1/2)^6`

`= "^6C_1 xx 1/64`        ...`[because ""^nC_x = "^nC_(n - x)]`

`= 6/64`

= `3/32`

Hence, the probability of 5 successes is `3/32`

shaalaa.com
  Is there an error in this question or solution?
Chapter 8: Binomial Distribution - Exercise 8.1 [Page 251]

Video TutorialsVIEW ALL [1]

RELATED QUESTIONS

A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of at least 5 successes.


A die is thrown 6 times. If ‘getting an odd number’ is a success, find the probability of at most 5 successes.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards; find the probability that all the five cards are spades.


Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards; find the probability that none is a spade.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that none of the floppy disc work.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly one floppy disc work.


In a box of floppy discs, it is known that 95% will work. A sample of three of the discs is selected at random. Find the probability that exactly two floppy disc work.


Choose the correct option from the given alternatives:

The mean and the variance of a binomial distribution are 4 and 2 respectively. Then the probability of 2 successes is


Choose the correct option from the given alternatives:

For a binomial distribution, n = 5. If P(X = 4) = P(X = 3), then p = ______


For a binomial distribution, n = 4. If 2P(X = 3) = 3P(X = 2), then p = ______.


If X ~ B(4, p) and P(X = 0) = `16/81`, then P(X = 4) = ______.


Let X ~ B(10, 0.2). Find P(X ≥ 1).


Let X ~ B(10, 0.2). Find P(X ≤ 8).


The probability that a bomb will hit a target is 0.8. Find the probability that out of 10 bombs dropped, exactly 2 will miss the target.


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: exactly one has a burst tyre


The probability that a mountain-bike travelling along a certain track will have a tyre burst is 0.05. Find the probability that among 17 riders: two or more have burst tyre.


The probability that a lamp in a classroom will be burnt out is 0.3. Six such lamps are fitted in the class-room. If it is known that the classroom is unusable if the number of lamps burning in it is less than four, find the probability that the classroom cannot be used on a random occasion.


A large chain retailer purchases a certain kind of electronic device from a manufacturer. The manufacturer indicates that the defective rate of the device is 3%. The inspector of the retailer picks 20 items from a shipment. What is the probability that the store will receive at most one defective item?


An examination consists of 10 multiple choice questions, in each of which a candidate has to deduce which one of five suggested answers is correct. A completely unprepared student guesses each answer completely randomly. What is the probability that this student gets 8 or more questions correct? Draw the appropriate morals.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that all 8 machines.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that 7 or 8 machines.


The probability that a machine will produce all bolts in a production run within specification is 0.998. A sample of 8 machines is taken at random. Calculate the probability that at most 6 machines will produce all bolts within specification. 


The probability that a machine develops a fault within the first 3 years of use is 0.003. If 40 machines are selected at random, calculate the probability that 38 or more will not develop any faults within the first 3 years of use.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 1.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 2.


A computer installation has 10 terminals. Independently, the probability that any one terminal will require attention during a week is 0.1. Find the probabilities that 3 or more, terminals will require attention during the next week.


It is observed that it rains on 12 days out of 30 days. Find the probability that it rains exactly 3 days of week.


If the probability of success in a single trial is 0.01. How many trials are required in order to have a probability greater than 0.5 of getting at least one success?


If E(x) > Var(x) then X follows _______.


Fill in the blank :

In Binomial distribution probability of success Remains constant / independent from trial to trial.


In Binomial distribution if n is very large and probability success of p is very small such that np = m (constant) then _______ distribution is applied.


If X ~ B(n, p) with n = 10, p = 0.4, then find E(X2).


In Binomial distribution, probability of success ______ from trial to trial


State whether the following statement is True or False:

For the Binomial distribution, Mean E(X) = m and Variance = Var(X) = m


In a binomial distribution `B(n, p = 1/4)`, if the probability of at least one success is greater than or equal to `9/10`, then n is greater than ______.


In a binomial distribution, n = 4 and 2P(X = 3) = 3P(X = 2), then q = ______.


A pair of dice is thrown 3 times. If getting a doublet is considered a success, find the probability of getting at least two success.

Solution:

A pair of dice is thrown 3 times.

∴ n = 3

Let x = number of success (doublets)

p = probability of success (doublets)

∴  p = `square`, q = `square`

∴ x ∼ B (n, p)

P(x) = nCxpx qn–x

Probability of getting at least two success means x ≥ 2.

∴ P(x ≥ 2) = P(x = 2) + P(x = 3)

= `square` + `square`

= `2/27`


If X is a binomial variable with range {0, 1, 2, 3, 4} and P(X = 3) = 3P(X = 4) then the parameter ‘p’ of the binomial distribution is


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×